OpenStudy (anonymous):

A 170lb baseball player slides into 2nd base. the player was running 8.2 m/s and traveled 5.5 ft. What is the coefficient of kinetic friction between the player and the ground while he is sliding?

OpenStudy (anonymous):

:)

OpenStudy (espex):

You need to start by putting all your data in a common format. Use the equation \[\sum F=F_{slide} - F_{fric}\] F=ma, friction=k*natural force, natural force=mg

OpenStudy (anonymous):

so the natural force would be 16.5?

OpenStudy (anonymous):

What does the k stand for?

OpenStudy (espex):

Frictional coefficient.

OpenStudy (anonymous):

Would that be his weight in newtons?

OpenStudy (espex):

A newton is defined as kg*m/s^2

OpenStudy (anonymous):

Yes, so would I just transfer 170 lb to newtons?

OpenStudy (anonymous):

Which would be 757.3 n

OpenStudy (espex):

\[170 lbs \times \frac{1kg}{2.2046lbs}=77.11kg\] \[77.11kg \times 9.8\frac{m}{s^2}=755.7N\]

OpenStudy (anonymous):

Okay would the final equation be \[755.7N \div 1.68g(9.8m/s ^{2}\]

OpenStudy (espex):

No, I don't know where you got 1.68g from.

OpenStudy (anonymous):

Oops nevermind, Im confused then on what to put in for the grams :/

OpenStudy (anonymous):

or the mass

OpenStudy (espex):

mass is kg, force is newtons, velocity is m/s, acceleration is m/s^2 assuming you are using metric.

OpenStudy (anonymous):

Yes I am using metric but how do I fill in the mass if I dont have it?

OpenStudy (espex):

You do have the mass.

OpenStudy (anonymous):

So I can use the baseball players weight for the newton and kg?

OpenStudy (espex):

The baseball player's "weight" is just his mass times the force of gravity.

OpenStudy (anonymous):

So will the equation be \[755.7N \div 76.9kg(9.8m/s ^{2)}\]

OpenStudy (espex):

The sum of the forces will be the force of the slide - the force of friction, force is equal to mass times acceleration, so you will have something like sumF=ma - mu x Fn and then solve for mu.

OpenStudy (espex):

If you draw a free body diagram you can see more clearly the forces at work.

OpenStudy (anonymous):

Hmm okay, Im just confused on what the force of friction is and the force of slide

OpenStudy (espex):

The force of friction is a force that causes an object to want to stay at rest. So if the ball player is sliding then the force of friction is what "pulls" the opposite direction to slow him down.

OpenStudy (anonymous):

Okay but how would I find that

OpenStudy (espex):

Okay, sorry for the delay, real life called. So our knowns \[v=8.2\frac{m}{s}, \space m=77.18kg, \space d=1.68m\] Since we have velocity and distance we can calculate time\[v=\frac{d}{t} \rightarrow t=\frac{d}{v} \rightarrow \frac{1.68m}{8.2\frac{m}{s^2}}=0.20s\] Now to find the force of friction, f=ma, we need to find his acceleration in the x direction (slowing down). \[a=\frac{v_f^2-v_i^2}{2\times (x_f-x_i)} \rightarrow \space \frac{0^2-8.2m^2}{2\times(1.68m-0)}=-20.38\] So the force of friction is\[F_{fric}=ma \rightarrow\space 77.18kg \times -20.38\frac{m}{s^2}=-1572.6N\] And you know that \[F_{fric}=\mu F_{nat} \rightarrow\space \mu mg\]You are looking for mu so solve algebraically and get\[\frac{F_{fric}}{mg}=\mu \rightarrow\space \frac{1572.6N}{77.15kg \times 9.8\frac{m}{s^2}}=2.079\]

OpenStudy (anonymous):

ok thank you so much :)